We can't figure out what the question is asking for. Please help. Here is the question.
Which of the following shows the right side of the equation after applying the reasonable step 2/3 -h/2=5/8
The answers are:
a)5/8+5/5
b) 5/8-5/8
c) 5/8+2/3
d) 5/8-2/3

Answer by richwmiller(17219)

Assume you are trying to solve the equation for h.
What is the next step?
you would subtract 2/3 from both sides.
2/3 -h/2=5/8
-h/2=5/8-2/3
so the answer is the last one
d) 5/8-2/3